# Dataset Condensation with Color Compensation

### v2

@@ 4.5 Training diffusion models with DC3
Figure 8: Fine-tuning diffusion models with DC3. To validate the effectiveness of our method, each fine-tuned model generates 10,000 images for efficient calculation according to the number of samples in ImageNette (12894).
Besides being effective on downstream tasks, we argue that the condensed datasets should adapt to pre-training or fine-tuning tasks. Based on our computational resources, we fine-tune the Stable Diffusion (SD) and Diffusion Transformers (DiT) (Peebles & Xie, 2023) on the original and mixed (original and DC3-condensed IPC-50 data) ImageNette dataset with LoRA (Hu et al., 2022) and DiffFit (Xie et al., 2023).
The results drawn in fig.˜8 demonstrate that datasets generated by DC3 effectively support the fine-tuning pipelines without inducing model collapse. The visualizations are in the supplementary materials. Specifically, the models fine-tuned by mixed data achieve better FID-IS scores compared to those on the original dataset and non-finetuned baselines, indicating better alignment with real data distributions. This performance gap stems from its ability to preserve feature diversity by chromatic alignment and submodular sampling, which maintains essential data manifolds while removing redundant information.
@@ Appendix A Experimental Details
The collaboration between sampling and compensation endows DC3 with powerful condensation capabilities. There are details of these pipelines that need to be explained. In submodular sampling, appendix˜B presents the results for different cluster numbers (K), as K significantly affects the sampling results. Also, in the color compensation phase, appendix˜C introduces the selection guideline of hue prompts. Additionally, appendix˜D experiments on different crop-and-stitch methods to further enhance the information density for the condensed images.
During the validation stage, the parameter details of all experiments are listed in table˜12. Finally, we provide a wealth of visualizations in appendices˜E, G and H, including condensed images and fine-tuned DiT-generated images, which prove the effectiveness of DC3.
Appendix B Selection of Cluster Numbers
Appendix C Guideline of Hue Prompt Selection
Cool Hue Warm Hue
rainy sepia
snowy sunny
infrared daylight
underwater vivid colors
frozen lake golden hour
Table 11: Examples of instruction prompts for cool and warm hues.
DC3 tried to execute the pixel-level optimization to mitigate color homogenization with instruction-conditioned diffusion models. The instructions are grouped by cool and warm hues. This taxonomy originates from quantitative analysis of natural lighting in photographic aesthetics: cool-hue scenes exhibit spectral energy concentrated in short wavelengths (400-500 nm), whereas warm hues correlate with long-wavelength distributions (550-700 nm). As listed in table˜11, cool-hue prompts (e.g., “rainy”, “snowy”) align with low-color-temperature scenes (e.g., blue-cyan hue in rain/snow). In contrast, warm-hue prompts (e.g., “sunny”, “golden hour”) capture high-color-temperature spectral characteristics (e.g., orange-yellow hue under strong illumination). By embedding these physically grounded semantic instructions into latent diffusion models, we successfully implement pixel-level color compensation (e.g., adjusting illumination while preserving structural integrity).
The selection strategy adheres to the following criteria: (a) visual discriminability — terms must induce color shifts aligned with human perception (e.g., “underwater” implies blue-green hue, “vivid colors” triggers saturation enhancement), (b) complementary coverage — prompts should span opposing color families to maximize gamut diversity.
@@ Appendix E Visualization Analysis
As a performance matching method, the KDE curves of MTT are different from other matching methods. From fig.˜10(a), the distribution reveals a leptokurtic profile characterized by a peaked central tendency and attenuated tails. This statistical property indicates significant Color Redundancy, where the compressed dataset over-represents narrow chromatic ranges while under-sampling critical color variations. Both Color Redundancy and Color Homogenization phenomena compressed data manifolds inadequately span the original color space, thereby violating the manifold hypothesis essential for dataset generalization. In contrast, DC3 fits the color distribution well on CIFAR-10 (fig.˜10(b)) and ImageNette (fig.˜10(c)) and achieves superior performance.
@@ Appendix G Visualizations from Fine-tuned DiT
Figure 11: Comparison of images generated by ImageNette fine-tuned DiT with different settings. The first row visualizes the results from the model fine-tuned with the original data, while the second row presents outputs from the mixed data (original ImageNette and DC3 condensed ImageNette IPC-50) fine-tuned DiT.

### v3

@@ 4.5 Training diffusion models with DC3
Figure 8: Fine-tuning diffusion models with DC3. To validate the effectiveness of our method, each fine-tuned model generates 10,000 images according to the number of samples in ImageNette (12894).
Besides being effective on downstream tasks, we argue that the condensed datasets should adapt to generation tasks. We fine-tune the Stable Diffusion (SD) and Diffusion Transformers (DiT) (Peebles & Xie, 2023) on the original and mixed (original and DC3-condensed IPC-50 data) ImageNette dataset with LoRA (Hu et al., 2022) and DiffFit (Xie et al., 2023).
The results drawn in fig.˜8 show that datasets generated by DC3 effectively support the fine-tune pipelines without model collapse. Specifically, the models fine-tuned by mixed data achieve better FID-IS (Salimans et al., 2016; Heusel et al., 2017) scores compared to those on the original dataset and baselines, indicating better alignment with real data distributions. This performance gap stems from its ability to preserve feature diversity by chromatic alignment and submodular sampling, which maintains essential data manifolds while removing redundant information.
@@ 5 Conclusion
Acknowledgments
This research was supported in part by JSPS KAKENHI Grant Numbers JP24K23849 and JP25K21218.
@@ Appendix A Experimental Details
The collaboration between sampling and compensation enables DC3 to possess powerful condensation capabilities. There are details of these pipelines that require explanation. In submodular sampling, appendix˜B presents the results for different cluster numbers (K), as K significantly affects the sampling results. Also, in the color compensation phase, appendix˜C introduces the selection guideline of hue prompts. Additionally, appendix˜D experiments on different crop-and-stitch methods to further enhance the information density for the condensed images.
During the validation stage, the parameter details of all experiments are listed in table˜11. Finally, we provide a wealth of visualizations in appendices˜E, G and H, including condensed images and fine-tuned DiT-generated images, which prove the effectiveness of DC3. For DC3 and the reproduced baselines, we report the mean and standard deviation over three runs. For other competing methods, we quote the results directly from their original publications.
@@
Appendix B Selection of Cluster Numbers
Appendix C Guideline of Hue Prompt Selection
Cool Hue Warm Hue
rainy sepia
snowy sunny
infrared daylight
underwater vivid colors
frozen lake golden hour
Table 12: Examples of instruction prompts for cool and warm hues.
DC3 attempted to implement pixel-level optimization to mitigate color homogenization using instruction-conditioned diffusion models. Cool and warm hues group the instructions. This taxonomy originates from quantitative analysis of natural lighting in photographic aesthetics: cool-hue scenes exhibit spectral energy concentrated in short wavelengths (400-500 nm), whereas warm hues correlate with long-wavelength distributions (550-700 nm). As listed in table˜12, cool-hue prompts (e.g., “rainy”, “snowy”) align with low-color-temperature scenes (e.g., blue-cyan hue in rain/snow). In contrast, warm-hue prompts (e.g., “sunny”, “golden hour”) capture high-color-temperature spectral characteristics (e.g., orange-yellow hue under strong illumination). By embedding these physically grounded semantic instructions into latent diffusion models, we successfully implement pixel-level color compensation (e.g., adjusting illumination while preserving structural integrity).
The selection strategy adheres to the following criteria: (a) visual discriminability — terms must induce color shifts aligned with human perception (e.g., “underwater” implies blue-green hue, “vivid colors” triggers saturation enhancement), (b) complementary coverage — prompts should span opposing color families to maximize gamut diversity. In our experiment, we randomly selected one prompt from both cool and warm hues for color compensation. Then we crop-and-stitch the generated image to create a single distilled image.
@@ Appendix E Visualization Analysis
As a performance matching method, the KDE curves of MTT are different from those of other matching methods. From fig.˜10(a), the distribution reveals a leptokurtic profile characterized by a peaked central tendency and attenuated tails. This statistical property indicates significant Color Redundancy, where the compressed dataset over-represents narrow chromatic ranges while under-sampling critical color variations. Both Color Redundancy and Color Homogenization phenomena compressed data manifolds inadequately span the original color space, thereby violating the manifold hypothesis essential for dataset generalization. In contrast, DC3 fits the color distribution well on CIFAR-10 (fig.˜10(b)) and ImageNette (fig.˜10(c)) and achieves superior performance.
@@ Appendix G Visualizations from Fine-tuned DiT
Figure 11: Comparison of images generated by ImageNette fine-tuned DiT with different settings. The first row displays outputs from DiT without fine-tuning. The second row visualizes the results from the model fine-tuned with the original data, while the third row presents outputs from the mixed data (original ImageNette and DC3 condensed ImageNette IPC-50) fine-tuned DiT.